# Complementary polynomials in quantum signal processing

### v1

@@ 1.5 Plan of the paper
Theorem 2 and Corollaries 2.1 and 2.2 are proved in Section 2. In Section 3, we design two algorithms for the computation of $Q$ following from Corollary 2.2. Error analysis of the algorithms is performed in Section 4 and numerical results comparing our algorithm to the optimization-based approach of [3] are presented in Section 5. Section 6 contains a discussion of our results and possibilities for future work. In Appendix A, we show that different QSP parameterizations can be viewed as special cases of GQSP.
@@ 2.1 Proof of Theorem 2
By construction, $U(z)$ is analytic on $\overline{\mathbb{D}}$. The real part of $U(z)$ is found to be
@@
The second case of (1.4c) is obtained from the first using the Plemelj formula [19]. Note that the integrand in (1.4a) has a simple pole at $z^{\prime}=z$ with residue $2\log\big{(}\frac{1-\lvert P(z)\rvert^{2}}{\lvert Q_{0}(z)\rvert^{2}}\big{)}$. Thus, applying the Plemelj formula as $z\in\mathbb{D}$ approaches the contour $\mathbb{T}$ gives (1.4b).
@@
$H[\mathrm{e}^{\mathrm{i}n\theta}]=\begin{cases}\mathrm{i}\mathrm{e}^{\mathrm{i}n\theta}&n\in\mathbb{Z}_{>0}\\ 0&n=0\\ -\mathrm{i}\mathrm{e}^{\mathrm{i}n\theta}&n\in\mathbb{Z}_{<0}.\end{cases}$ (2.12)
@@ 3 Numerical methods
To numerically evaluate (3.3), we make two approximations that allow us to compute $Q$ in the monomial basis by a sequence of FFTs. As the performance of an FFT is optimized when the number of Fourier modes is a power of $2$, we choose the size of this basis to be $N=2^{M}$ for some $M\in\mathbb{Z}_{\geq 1}$ satisfying $M\geq\lceil\log_{2}(d+1)\rceil$. The error analysis of Algorithm 1, which will result from the approximations and analysis in this section, is performed in Section 4. First, we introduce the Laurent polynomial truncation of (3.1)
@@
Second, we will approximate the coefficients (3.2) by discrete Fourier transforms. Consider the root of unity
$\omega_{N}\coloneqq\mathrm{e}^{2\pi\mathrm{i}/N}.$ (3.5)
@@ 3.1 Algorithm
We combine the observations obtained in this section into the following algorithms to compute $\tilde{Q}_{2,N}(z)$ (3.13), the numerical approximation to the canonical complementary polynomial to $P(z)$. In order to avoid costly root-finding to determine $Q_{0}(z)$, we will consider situations where $Q_{0}(z)=1$, see Algorithm 1. However, this is not a restriction; by scaling down the input polynomial ever so slightly, $Q_{0}(z)=1$ can always be achieved, leading to the general Algorithm 2 applicable to any polynomial.
For many practical applications of QSP-type algorithms in quantum computation, a parameter $\delta>0$ with
$\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 1-\delta$ (3.14)
can be controlled a priori in the construction of the polynomial $P(z)$ approximating a target function. Then, (3.14) ensures that $1-|P(z)|^{2}$ has no roots on the unit circle $z\in\mathbb{T}$, and $Q_{0}(z)=1$. The numerical method explained above can then be used directly, and is summarised in the following Algorithm 1:
Algorithm 1 (Construction of the complementary polynomial for known $\delta$)
@@
  • The coefficients (in the monomial basis) of a polynomial $P\in\mathbb{C}[z]$ of degree $d$ with $\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 1-\delta$, for a $\delta>0$.
  • An integer $N\geq d$ controlling the accuracy $\varepsilon$ of the output, called dimension of the FFT.
@@
  • The coefficients (in the monomial basis) of the canonical complementary polynomial $Q\in\mathbb{C}[z]$ of degree $d$ from Theorem 2, up to accuracy $\varepsilon$.
@@
  • Runtime: $O(N\log N)$
  • Sufficient $N$ for accuracy $\varepsilon$: $N=O\!\left(\frac{d}{\delta}\log\frac{d}{\delta\varepsilon}\right)$, see Theorem 3
@@
    Compute $(P(\omega_{N}^{n}))_{n=0}^{N-1}$, the input polynomial $P(z)=\sum_{n=0}^{d}p_{n}z^{n}$ evaluated at all $N$th roots of unity, with the inverse FFT of $(p_{n})_{n=0}^{N-1}$ padded with zeros, i.e. $p_{n}=0,n>d$.
@@
  • See Figure 1 for Python code, and Section 3 for numerical results.
The algorithm relies heavily on the fast Fourier transform (FFT) to switch between coefficients of polynomials and values of polynomials at roots of unity, and to apply the Fourier multiplier in Fourier space. The FFTs of dimension $N$ result in the runtime $O(N\log N)$.
In Theorem 3 in the next section, we prove that Algorithm 1 computes the canonical complementary polynomial (the unique solution, up to phase, with no roots in $\mathbb{D}$, as in Theorem 2), up to accuracy $\varepsilon$ in the coefficients. Moreover, the same theorem shows that the algorithm is efficient in degree and error because of its asymptotic scaling $N=O(d\log(d/\varepsilon))$.
When $\delta$ is very small, the scaling $1/\delta$ from the bound on $N$ in Theorem 3 suggests a long runtime; in the extreme case $\delta=0$, the proof for Algorithm 1 fails because $Q_{0}(z)\neq 1$. For those cases, we present Algorithm 2, in which the initial polynomial is first scaled down as $P(z)\to(1-\tfrac{\varepsilon}{4})P(z)$ to achieve an effective $\delta=\varepsilon/4$. We will prove in Theorem 4 that the polynomial generated by Algorithm 2 with asymptotic scaling $N=O(d/\varepsilon\,\log(d/\varepsilon))$ satisfies the complementarity condition up to accuracy $\varepsilon$, i.e.,
$\big{\lVert}\lvert P(z)\rvert^{2}+\lvert Q(z)\rvert^{2}-1\big{\rVert}_{\infty,\mathbb{T}}\leq\varepsilon.$ (3.15)
Closeness in complementarity condition is a weaker statement than closeness to the exact canonical complementary polynomial like the output of Algorithm 1. Yet, it enables us to rigorously and efficiently extend our numerical method to all polynomials:
Algorithm 2 (Construction of the complementary polynomial for unknown, small, or zero $\delta$)
@@
  • The coefficients (in the monomial basis) of a polynomial $P\in\mathbb{C}[z]$ of degree $d$ with $|P(z)|\leq 1$, $z\in\mathbb{T}$.
  • An integer $N\geq d$ controlling the accuracy $\varepsilon$ of the output, called dimension of the FFT.
@@
  • The coefficients (in the monomial basis) of a polynomial $Q\in\mathbb{C}[z]$ of degree $d$ that satisfies the complementarity condition up to accuracy $\varepsilon$, i.e., $\big{\lVert}|P(z)|^{2}+|Q(z)|^{2}-1\big{\rVert}_{\infty,\mathbb{T}}\leq\varepsilon$.
@@
  • Runtime: $O(N\log N)$
  • Sufficient $N$ for accuracy $\varepsilon$: $N=O\!\left(\frac{d}{\varepsilon}\log\frac{d}{\varepsilon}\right)$, see Theorem 4
@@
    Return the result of Algorithm 1 with input the downscaled polynomial from the previous step, for which $\delta=\varepsilon/4$, and $N$ chosen to yield an accuracy of $\varepsilon/6(d+1)$.
@@
Empirically, we find that Algorithm 2 may not be needed. Even without the initial downscaling, our numerical results in Section 5 suggest that Algorithm 1 alone is efficacious even when $\delta=0$, with error dependence $N\approx O(d/\sqrt[4]{\varepsilon})$. Here, and in our error analysis in the next section, we prefer to include Algorithm 2 to preserve full mathematical rigor.
@@ 4 Error analysis of algorithms
We perform error analysis on Algorithm 1 and Algorithm 2, developed in the previous section. We state our main results, Theorem 3, Corollary 3.1, and Theorem 4 below. The corresponding proofs are given in Sections 4.3, 4.4, and 4.5, respectively.
@@ 4.2 Results of error analysis
Here, we establish rigorous error bounds on Algorithms 1 and 2, using the error metrics introduced in the previous subection.
@@
Suppose $P\in\mathbb{C}[z]$ satisfying $\lVert P(z)\rVert_{\mathbb{T},\infty}\leq 1-\delta$ for some $\delta\in(0,1)$ and $\varepsilon\in(0,1)$ are given. Choose $N\in\mathbb{Z}_{\geq 1}$ such that
@@
Letting $\varepsilon^{\prime}=\tilde{\Phi}(P,\tilde{Q}_{2,N})$, we see from Theorem 3 that for fixed $\delta$, $N=O(d\log(1/\varepsilon^{\prime}))$ is required to achieve (4.13). Numerical results verifying this assertion are presented in Section 5.
@@ 4.3 Proof of Theorem 3
$M(r)\coloneqq\max_{\rho=1/r,r}\max_{z\in\mathbb{T}}\big{\lvert}\log\big{(}{1-P(\rho z)P^{*}(1/\rho z)}\big{)}\big{\rvert}\quad(r\in(1,R)).$ (4.18)
@@
To obtain a similar bound for the difference between $\Pi[S_{N}(\mathrm{e}^{\mathrm{i}\theta})]$ and $\Pi[\tilde{S}_{N}(\mathrm{e}^{\mathrm{i}\theta})]$, we recall the discrete Poisson summation formula [23, Chapter 6]
@@
Next, we analyze the asymptotic behaviour of our bound (4.8) on a sufficient $N$. For this we first introduce the parameter
@@
$N_{0}(\varepsilon,\delta,d)\sim 2dc_{\delta}\log\frac{1}{\varepsilon}=O\!\left(\log\frac{1}{\varepsilon}\right),\quad(\varepsilon\to 0,\delta,d\ \text{fixed})$ (4.49)
@@
Next, note that $c_{\delta}d\to+\infty$ as $\delta\to 0^{+}$ and/or $d\to\infty$, such that we have the asymptotic equation
$\frac{1}{r_{\delta}-1}=\frac{1}{\exp(\frac{1}{c_{\delta}d})-1}\sim c_{\delta}d\quad(\delta\to 0\ \text{and/or}\ d\to\infty).$ (4.50)
We insert this into (4.48), but keep $\varepsilon$ because we will take a limit $\varepsilon\to 0$ later:
$N_{0}(\varepsilon,\delta,d)$ $\sim 2dc_{\delta}\log\frac{1}{\varepsilon}+2dc_{\delta}\log\left(8\log\frac{1}{\delta}\right)+2dc_{\delta}\log\left(c_{\delta}d\right)\quad(\delta\to 0\ \text{and/or}\ d\to\infty).$ (4.51)
The middle term is subdominant: In the case $d\to\infty$, the middle logarithm is a constant, and in the case $\delta\to 0$ note that $c_{\delta}\sim 1/\delta$. Dropping the middle term results in
$N_{0}(\varepsilon,\delta,d)\sim 2dc_{\delta}\log\left(\frac{c_{\delta}d}{\varepsilon}\right)\quad(\delta\to 0\ \text{and/or}\ d\to\infty).$ (4.52)
@@
$N_{0}(\varepsilon,\delta,d)\sim 2dc_{\delta}\log\frac{d}{\varepsilon}=O\!\left(d\log\frac{d}{\varepsilon}\right)\quad(\varepsilon\to 0,d\to\infty,\delta\ \text{fixed}).$ (4.53)
As $\delta\to 0$, we get
$N_{0}(\varepsilon,\delta,d)\sim 2\frac{d}{\delta}\log\left(\frac{d}{\varepsilon\delta}\right)\quad(\delta\to 0).$ (4.54)
which is valid regardless of whether $\varepsilon\to 0$ and/or $d\to\infty$. In particular, (4.11) holds provided at least one of $\varepsilon\to 0$, $\delta\to 0$, or $d\to\infty$.
@@
$\lVert Q(z)-\tilde{Q}_{2,N}(z)\rVert_{\mathbb{T},\infty}<(d+1)\varepsilon$ (4.55)
@@
$\big{\lVert}|Q(z)|^{2}-|\tilde{Q}_{2,N}(z)|^{2}\big{\rVert}_{\infty,\mathbb{T}}=$ $\;\lVert 2Q^{\prime}(z)-Q^{\prime}(z)+\tilde{Q}(z)\rVert_{\infty,\mathbb{T}}\cdot\big{\lVert}Q(z)-\tilde{Q}_{2,N}(z)\big{\rVert}_{\infty,\mathbb{T}}$
$<$ $\;(2+(d+1)\epsilon)(d+1)\varepsilon)<(d+1)(d+3)\varepsilon.$ (4.56)

### v2

@@ 1.5 Plan of the paper
Theorem 2 and Corollaries 2.1 and 2.2 are proved in Section 2. In Section 3, we design two algorithms for the computation of $Q$ following from Corollary 2.2. Error analysis of our algorithms is performed in Section 4 and numerical results comparing our algorithm to the optimization-based approach of [3] are presented in Section 5. Section 6 contains a discussion of our results and possibilities for future work. In Appendix A, we show that different QSP parameterizations can be viewed as special cases of GQSP. Appendix B contains precise statements of the complex analysis theorems used to prove our results.
@@ 2.1 Proof of Theorem 2
By construction, $U(z)$ is holomorphic on $\overline{\mathbb{D}}$. The real part of $U(z)$ is found to be
@@
The second case of (1.4c) is obtained from the first using the Plemelj formula (B.5). Note that the integrand in (1.4a) has a simple pole at $z^{\prime}=z$ with residue $2\log\big{(}\frac{1-\lvert P(z)\rvert^{2}}{\lvert Q_{0}(z)\rvert^{2}}\big{)}$. Thus, applying the Plemelj formula as $z\in\mathbb{D}$ approaches the contour $\mathbb{T}$ gives (1.4b).
@@
$H[\mathrm{e}^{\mathrm{i}n\theta}]=\begin{cases}\mathrm{i}\mathrm{e}^{\mathrm{i}n\theta}&n\in\mathbb{Z}_{\geq 1}\\ 0&n=0\\ -\mathrm{i}\mathrm{e}^{\mathrm{i}n\theta}&n\in\mathbb{Z}_{\geq 1}.\end{cases}$ (2.12)
@@ 3 Numerical methods
To numerically evaluate (3.3), we make two approximations that allow us to compute $Q$ in the monomial basis by a sequence of FFTs. As the performance of an FFT is optimized when the number of Fourier modes is a power of $2$, we choose the size of this basis to be $N=2^{M}$ for some $M\in\mathbb{Z}_{\geq 1}$ satisfying $M\geq\lceil\log_{2}(d+1)\rceil$. The error analysis of Algorithm 1, which will result from the approximations and analysis in this section, is performed in Section 4. First, we introduce the Laurent polynomial truncation of (3.1),
@@
Second, we will approximate the coefficients (3.2) by discrete Fourier transforms. Consider the primitive $N$th root of unity
$\omega_{N}\coloneqq\mathrm{e}^{2\pi\mathrm{i}/N},$ (3.5)
@@ 3.1 Algorithm
We combine the observations obtained in this section into algorithms to compute $\tilde{Q}_{2,N}$ (3.13), an approximate canonical complementary polynomial to $P$. In order to avoid the use of root-finding to determine $Q_{0}$, we will consider situations where $Q_{0}(z)=1$; see Algorithm 1. However, this is not a restriction; by downscaling the input polynomial, $Q_{0}(z)=1$ can always be achieved. Accordingly, the generalized Algorithm 2 applies to any target polynomial satisfying $\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 1$.
For many practical applications of QSP-type algorithms in quantum computation, the parameter $\delta$ in (1.9) can be controlled a priori in the construction of a polynomial $P$ approximating a target function. Then, (1.9) ensures that $1-|P(z)|^{2}$ has no roots on $\mathbb{T}$ and hence $Q_{0}(z)=1$. Algorithm 1 applies directly in this scenario.
Algorithm 1 (Construction of a canonical complementary polynomial for known $\delta$)
@@
  • The monomial coefficients $(p_{n})_{n=0}^{d}$ of $P\in\mathbb{C}[z]$, $\deg P=d$, satisfying (1.9) for known $\delta\in(0,1)$.
  • An integer $N\in\mathbb{Z}_{\geq d}$ determining the dimension of the FFTs and thus controlling the accuracy $\varepsilon$ of the output.
@@
  • The monomial coefficients $(\tilde{q}_{n})_{n=0}^{d}$ of an approximate canonical complementary polynomial $\tilde{Q}_{2,N}\in\mathbb{C}[z]$, $\deg\tilde{Q}_{2,N}=d$, approximating $Q$ from Theorem 2 to accuracy $\varepsilon$ in each monomial coefficient; see (3.13).
@@
  • Runtime: $O(N\log N)$.
  • Sufficient $N$ for accuracy $\varepsilon$: $N=O\!\left(\frac{d}{\delta}\log\frac{d}{\delta\varepsilon}\right)$; see Theorem 3.
@@
    Compute $(P(\omega_{N}^{n}))_{n=0}^{N-1}$, the input polynomial $P(z)=\sum_{n=0}^{d}p_{n}z^{n}$ evaluated at all $N$th roots of unity, with the inverse FFT of $(p_{n})_{n=0}^{N-1}$ padded with zeros, i.e., $p_{n}=0$ for $n>d$.
@@
  • See Figure 1 for Python code and Section 3 for numerical results.
The algorithm relies on FFTs to map between the coefficients of a polynomial and its values at roots of unity and to apply the Fourier multiplier $\Pi$ in Fourier space (1.6)–(1.8). The FFTs of dimension $N$ correspond to an overall runtime of $O(N\log N)$.
In Theorem 3, stated in Section 4, we prove that Algorithm 1 computes a canonical complementary polynomial to accuracy $\varepsilon$ in the monomial coefficients. We emphasize that the canonical complementary polynomial is the unique solution of Problem 1, up to a multiplicative phase, with no roots in $\mathbb{D}$, as in Theorem 2. Theorem 3 moreover shows that the algorithm is efficient in degree and error, with a sufficient $N=O\big{(}d\log\frac{d}{\varepsilon}\big{)}$ for fixed $\delta$.
For small $\delta$, the scaling $N\sim\frac{1}{\delta}$ from Theorem 3 suggests a long runtime; in the extreme case $\delta=0$, the proof for Algorithm 1 fails because $Q_{0}(z)\neq 1$. For those cases, we present Algorithm 2, in which the initial polynomial is downscaled as $P(z)\to(1-\tfrac{\varepsilon}{4})P(z)$ to achieve an effective $\delta=\frac{\varepsilon}{4}$. In Theorem 4, stated in Section 4, we prove that the polynomial generated by Algorithm 2 with $N=O\big{(}\frac{d}{\varepsilon}\log\frac{d}{\varepsilon})$ satisfies the complementarity condition (1.1) to accuracy $\varepsilon$, i.e., $\big{\lVert}\lvert P(z)\rvert^{2}+\lvert Q(z)\rvert^{2}\big{\rVert}_{\infty,\mathbb{T}}<\varepsilon$. Closeness in the complementarity condition (1.1) is a weaker statement than closeness to an exact canonical complementary polynomial, as Theorem 3 promises for Algorithm 1. Yet, it enables us to rigorously and efficiently extend our numerical method to all polynomials $P$ with $\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 1$.
Algorithm 2 (Construction of a complementary polynomial for zero, unknown, or small $\delta$)
@@
  • The monomial coefficients of $P\in\mathbb{C}[z]$, $\deg P=d$, with $\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 1$.
  • An integer $N\in\mathbb{Z}_{\geq d}$ determining the dimension of the FFTs and thus controlling the accuracy $\varepsilon$ of the output.
@@
  • The monomial coefficients $(\tilde{q}_{n})_{n=0}^{d}$ of a canonical complementary polynomial $\tilde{Q}_{2,N}\in\mathbb{C}[z]$, $\deg\tilde{Q}_{2,N}=d$, to accuracy $\varepsilon$ in the complementarity condition (1.1), i.e., $\big{\lVert}\lvert P(z)\rvert^{2}+\lvert Q(z)\rvert^{2}\big{\rVert}_{\infty,\mathbb{T}}<\varepsilon$.
@@
  • Runtime: $O(N\log N)$.
  • Sufficient $N$ for accuracy $\varepsilon$: $N=O\!\left(\frac{d}{\varepsilon}\log\frac{d}{\varepsilon}\right)$; see Theorem 4.
@@
    Return the result of Algorithm 1 with input the downscaled polynomial from the previous step, for which $\delta=\frac{\varepsilon}{4}$, and $N$ chosen to yield an accuracy of $\frac{\varepsilon}{5(d+1)}$.
@@
Empirically, we find that Algorithm 2 may not be needed. Even without the initial downscaling, our numerical results in Section 5 suggest that Algorithm 1 alone is efficacious even when $\delta=0$, with a sufficient $N=O\big{(}\frac{d}{\sqrt[4]{\varepsilon}}\big{)}$. Here, and in our error analysis in the next section, we include Algorithm 2 to preserve full mathematical rigor.
@@ 4 Error analysis of algorithms
We perform error analysis on Algorithm 1 and Algorithm 2, developed in the previous section. Our main results, Theorem 3, Corollary 3.1, and Theorem 4, are stated below. The corresponding proofs are given in Sections 4.3, 4.4, and 4.5, respectively.
@@ 4.2 Results of error analysis
We establish rigorous error bounds on Algorithms 1 and 2, using the error metrics introduced in the previous subsection.
@@
Suppose $P\in\mathbb{C}[z]$ satisfying (1.9) for some $\delta\in(0,1)$ and $\varepsilon\in(0,1)$ are given. Choose $N\in\mathbb{Z}_{\geq 1}$ such that
@@
Letting $\tilde{\varepsilon}=\tilde{\Phi}(P,\tilde{Q}_{2,N})$, we see from Theorem 3 that for fixed $\delta$, $N=O\big{(}d\log\frac{1}{\tilde{\varepsilon}}\big{)}$ is required to achieve (4.13). Numerical results verifying this assertion are presented in Section 5.
@@ 4.3 Proof of Theorem 3
$M(r)\coloneqq\max_{\rho=\frac{1}{r},r}\max_{z\in\mathbb{T}}\big{\lvert}\log\big{(}{1-P(\rho z)P^{*}(1/\rho z)}\big{)}\big{\rvert}\quad(r\in(1,R)).$ (4.18)
@@
To obtain a corresponding bound for the difference between $\Pi[S_{N}(\mathrm{e}^{\mathrm{i}\theta})]$ and $\Pi[\tilde{S}_{N}(\mathrm{e}^{\mathrm{i}\theta})]$, we recall the discrete Poisson summation formula [22, Chapter 6]
@@
Next, we analyze the asymptotic behavior of our bound (4.8) on a sufficient $N$. To this end, we introduce the parameter
@@
$N_{0}(\varepsilon,\delta,d)\sim 2dc_{\delta}\log\frac{1}{\varepsilon}=O\!\left(\log\frac{1}{\varepsilon}\right),\quad(\varepsilon\downarrow 0,\delta,d\ \text{fixed})$ (4.49)
@@
Next, note that $c_{\delta}d\to+\infty$ as $\delta\downarrow 0$ or $d\to\infty$, such that we have the asymptotic equation
$\frac{1}{r_{\delta}-1}=\frac{1}{\exp(\frac{1}{c_{\delta}d})-1}\sim c_{\delta}d\quad(\delta\downarrow 0\ \text{or}\ d\to\infty).$ (4.50)
We insert this into (4.48), but keep $\varepsilon$ because we will take a limit $\varepsilon\downarrow 0$ later:
$N_{0}(\varepsilon,\delta,d)$ $\sim 2dc_{\delta}\log\frac{1}{\varepsilon}+2dc_{\delta}\log\left(8\log\frac{1}{\delta}\right)+2dc_{\delta}\log\left(c_{\delta}d\right)\quad(\delta\downarrow 0\ \text{or}\ d\to\infty).$ (4.51)
The middle term is subdominant. In the case $d\to\infty$, the middle logarithm is a constant, and in the case $\delta\downarrow 0$ note that $c_{\delta}\sim 1/\delta$. Dropping the middle term results in
$N_{0}(\varepsilon,\delta,d)\sim 2dc_{\delta}\log\left(\frac{c_{\delta}d}{\varepsilon}\right)\quad(\delta\downarrow 0\ \text{or}\ d\to\infty).$ (4.52)
@@
$N_{0}(\varepsilon,\delta,d)\sim 2dc_{\delta}\log\frac{d}{\varepsilon}=O\!\left(d\log\frac{d}{\varepsilon}\right)\quad(\varepsilon\downarrow 0,d\to\infty,\delta\ \text{fixed}).$ (4.53)
As $\delta\downarrow 0$, we get
$N_{0}(\varepsilon,\delta,d)\sim 2\frac{d}{\delta}\log\left(\frac{d}{\varepsilon\delta}\right)\quad(\delta\downarrow 0).$ (4.54)
which is valid regardless of whether $\varepsilon\downarrow 0$ or $d\to\infty$. In particular, (4.11) holds provided $\varepsilon\downarrow 0$, $\delta\downarrow 0$, or $d\to\infty$.
@@
$\big{\lVert}Q(z)-\tilde{Q}_{2,N}(z)\big{\rVert}_{\mathbb{T},\infty}<(d+1)\varepsilon$ (4.55)
@@
$\big{\lVert}|Q(z)|^{2}-|\tilde{Q}_{2,N}(z)|^{2}\big{\rVert}_{\infty,\mathbb{T}}\leq$ $\;\big{\lVert}2Q(z)-Q(z)+\tilde{Q}_{2,N}(z)\big{\rVert}_{\infty,\mathbb{T}}\big{\lVert}Q(z)-\tilde{Q}_{2,N}(z)\big{\rVert}_{\infty,\mathbb{T}}$
$<$ $\;(2+(d+1)\epsilon)(d+1)\varepsilon<(d+1)(d+3)\varepsilon.$ (4.56)
It follows from (4.1) and (4.4) that